Decide how many solutions the equation has: y=1/3 x^2 +2x-3
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

y = (1/3)x^2 + 2x - 3

Discriminant = b^2 - 4ac = 4 - 4(1/3)(-3) = 8

The discriminant is greater than 0, so this equation has 2 real roots.

Answer by ikleyn(52754)

.

What is written in your post IS NOT an equation.

It is the definition of the function.


THEREFORE, your question is NONSENSICAL.


The post by @Penguin, who tries to answer nonsensical post, is NONSENSICAL itself, too.